# Structure-preserving approximation of the non-isothermal Cahn-Hilliard system

### v1

@@ 6 Numerical experiments
with parameters $a=0.01,\leavevmode\nobreak\ b=1,\leavevmode\nobreak\ d=1$ and $\vartheta_{c}=3$. The graphs are depicted in Figure 1. Mobility and heat conductivity are $\mathbf{M}=10^{-2}\cdot\mathbf{I}$ and $\mathbf{K}=5\cdot 10^{-2}\cdot\mathbf{I}$. The cross-coupling term $\mathbf{C}$ will be either $10^{-4}\cdot\mathbf{I}$ or $0\cdot\mathbf{I}$. The interface parameter is fixed to $\gamma=10^{-3}$. For simplicity we take $g^{*}$ in (10)–(12) are taken to be $g^{n}$. It means that they are evaluated explicitly and we set $\mu_{h}^{0}=0$. The resulting system is still nonlinear and solved using the Newton method with error tolerance $10^{-12}$. The resulting linear system is solved using a direct solver. The numerical method is implemented in NGSolve [36].
$\phi$$f(\phi,\vartheta)$$f(\phi,4)$
$\phi$$f(\phi,\vartheta)$$f(\phi,2)$
@@ 6.1 Convergence test
First we want to present experimental convergence results and start with the convergence in space. To this end, we consider the initial data
@@
$\operatorname{trans}_{[0,1]}(z)=\frac{\min\Big{(}\max\big{(}\tanh(z),-1\big{)},1\Big{)}+1}{2}\in[0,1].$
Since no exact solution is available, we will study self-convergence by computing the errors of the Cauchy sequence. Specifically, the error quantities we consider in space are given as
$\mathrm{err}(\nabla\phi,h_{k})$ $:=\max_{n\in\mathcal{I}_{\tau}}\|\phi^{n}_{h_{k}}-\phi^{n}_{h_{k+1}}\|_{1}^{2},\quad\mathrm{err}(\nabla\mu,h_{k}):=\tau\sum_{n=1}^{n_{\tau}}\|\mu^{n}_{h_{k}}-\mu^{n}_{h_{k+1}}\|_{1}^{2},$
$\mathrm{err}(\vartheta,h_{k})$ $:=\max_{n\in\mathcal{I}_{\tau}}\|\vartheta^{n}_{h_{k}}-\vartheta^{n}_{h_{k+1}}\|_{0}^{2},\quad\mathrm{err}(\nabla\vartheta,h_{k}):=\tau\sum_{n=1}^{n_{\tau}}\|\vartheta^{n}_{h_{k}}-\vartheta^{n}_{h_{k+1}}\|_{1}^{2}.$
@@
To this end, we consider mesh refinements with the following mesh sizes $h_{k}\approx 2^{-k}$ for $k=4,\ldots,8$ and for the time step we choose $\tau=2^{-10}\cdot 10^{-1}$. The final time is set to $T=10^{-1}$. The errors and approximate convergence rates, with $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, can be found in Table 1. They indicate optimal first-order convergence in the $H^{1}$-norm and the second-order convergence in the $L^{2}$-norm.
@@
4 $1.36\cdot 10^{-3}$ $8.37\cdot 10^{-5}$ $5.62\cdot 10^{-3}$ $3.88\cdot 10^{0}$
5 $4.35\cdot 10^{-4}$ $1.65$ $3.17\cdot 10^{-5}$ $1.40$ $2.96\cdot 10^{-4}$ $4.24$ $9.85\cdot 10^{-1}$ $1.98$
6 $1.25\cdot 10^{-4}$ $1.79$ $7.26\cdot 10^{-6}$ $2.13$ $1.90\cdot 10^{-5}$ $3.96$ $2.54\cdot 10^{-1}$ $1.96$
7 $3.21\cdot 10^{-5}$ $1.97$ $1.62\cdot 10^{-6}$ $2.16$ $1.27\cdot 10^{-6}$ $3.91$ $6.25\cdot 10^{-2}$ $2.02$
For the time convergence, the expected optimal order is first-order convergence. Therefore, we use time step refinements $\tau_{k}=2^{-k}\cdot 10^{-1}$ for $k=7,...,11$ and simulate again until a final time $T=10^{-1}$. The spatial mesh is fixed with mesh size $h=2^{-7}$. Here we consider similar errors as above, namely:
$\text{err}(\nabla\phi,\tau_{k})$ $:=\max_{n\in\mathcal{I}_{\tau_{k}}}\|\phi^{n}_{h,\tau_{k}}-\phi^{n}_{h,\tau_{k+1}}\|_{1}^{2},\leavevmode\nobreak\ \text{err}(\nabla\mu,\tau_{k}):=\tau_{k}\sum_{n=1}^{n_{\tau_{k}}}\|\mu^{n}_{h,\tau_{k}}-\mu^{n}_{h,\tau_{k+1}}\|_{1}^{2},$
$\text{err}(\vartheta,\tau_{k})$ $:=\max_{n\in\mathcal{I}_{\tau_{k}}}\|\vartheta^{n}_{h,\tau_{k}}-\vartheta^{n}_{h,\tau_{k+1}}\|_{0}^{2},\leavevmode\nobreak\ \text{err}(\nabla\vartheta,\tau_{k}):=\tau_{k}\sum_{n=1}^{n_{\tau_{k}}}\|\vartheta^{n}_{h,\tau_{k}}-\vartheta^{n}_{h,\tau_{k+1}}\|_{1}^{2}.$
@@
for $a\in\{\nabla\phi,\nabla\mu,\vartheta,\nabla\vartheta\}$. The errors and rates, for $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, can be found in Table 2. As expected the optimal first-order convergence in time is achieved.
@@
7 $2.08\cdot 10^{-7}$ $1.72$ $7.97\cdot 10^{-9}$ $1.94$ $4.43\cdot 10^{-8}$ $2.01$ $1.25\cdot 10^{-5}$ $2.01$
8 $5.89\cdot 10^{-8}$ $1.82$ $2.04\cdot 10^{-9}$ $1.97$ $1.11\cdot 10^{-8}$ $2.00$ $3.11\cdot 10^{-6}$ $2.01$
9 $1.59\cdot 10^{-8}$ $1.89$ $5.16\cdot 10^{-10}$ $1.98$ $2.76\cdot 10^{-9}$ $2.00$ $7.76\cdot 10^{-7}$ $2.00$
10 $4.15\cdot 10^{-9}$ $1.94$ $1.30\cdot 10^{-10}$ $1.99$ $6.90\cdot 10^{-10}$ $2.00$ $1.94\cdot 10^{-7}$ $2.00$
@@ 6.2 Illustrating example
Figure 3: Snapshots of volume fraction $\vartheta$ for different cross-coupling matrices $\mathbf{C}$: (Top) $\mathbf{C}=0\cdot\mathbf{I}$; (Bottom) $\mathbf{C}=10^{-4}\cdot\mathbf{I}$.
@@ 7 Conclusion & Outlook
In this work we have proposed a fully discrete finite element method for the non-isothermal Cahn-Hilliard equation based on the use of the entropy equation instead of the internal energy equation. The scheme preserves the conservation of mass and the entropy production, while the internal energy, here the same as total energy, is dissipated in time due to numerical diffusion. Numerical experiments document optimal order of convergence, i.e. the first order in time and the first or second order in space depending on the norm. We also illustrate the impact of the cross-coupling terms between the Cahn-Hilliard and the entropy equation by numerical simulations. In future work we want to extend our scheme to eliminate the numerical dissipation such that the discrete energy is conserved. Furthermore, we will consider extensions to incompressible flows.
Acknowledgement
The present research has been support by the German Science Foundation under the grant TRR146 Multiscale Simulation Methods for Soft Matter Systems and by the grant SPP 2256 Variational Methods for Predicting Complex Phenomena in Engineering Structures and Materials. M.L.-M. gratefully acknowledges the support of the Gutenberg Research Fellowship and the Mainz Institute for Multiscale Modelling (M3odel).

### v2

@@ 6 Numerical experiments
with parameters $a=0.01,\penalty 10000\ b=1,\penalty 10000\ d=1$ and $\vartheta_{c}=3$. The graphs are depicted in Figure 1. Mobility and heat conductivity are $\mathbf{M}=10^{-2}\cdot\mathbf{I}$ and $\mathbf{K}=5\cdot 10^{-3}\cdot\mathbf{I}$. The cross-coupling term $\mathbf{C}$ will be either $10^{-4}\cdot\mathbf{I}$ or $0\cdot\mathbf{I}$. The interface parameter is fixed to $\gamma=10^{-4}$. For simplicity we take $g^{*}$ in (14)–(16) are taken to be $g^{n}$. It means that they are evaluated explicitly and we set $\mu_{h}^{0}=0$. The resulting system is still nonlinear and solved using the Newton method with error tolerance $10^{-12}$. The resulting linear system is solved using a direct solver. The numerical method is implemented in NGSolve [36].
[…]
@@ 6.1 Convergence test
We now present experimental results to verify the spatial and temporal convergence of the proposed scheme. We consider the following initial data
@@
$\operatorname{trans}_{[0,1]}(z)=\frac{-\tanh(z)+1}{2}\in[0,1].$
As no exact solution is available, we will study self-convergence by computing the errors of the Cauchy sequence. We consider the following error norms:
$\mathrm{err}(\nabla\phi,h_{k})$ $:=\max_{n\in\mathcal{I}_{\tau}}\|\phi^{n}_{h_{k}}-\phi^{n}_{h_{k+1}}\|_{1},\quad\mathrm{err}(\nabla\mu,h_{k}):=\sqrt{\tau\sum_{n=1}^{n_{\tau}}\|\mu^{n}_{h_{k}}-\mu^{n}_{h_{k+1}}\|_{1}^{2}},$
$\mathrm{err}(\vartheta,h_{k})$ $:=\max_{n\in\mathcal{I}_{\tau}}\|\vartheta^{n}_{h_{k}}-\vartheta^{n}_{h_{k+1}}\|_{0},\quad\mathrm{err}(\nabla\vartheta,h_{k}):=\sqrt{\tau\sum_{n=1}^{n_{\tau}}\|\vartheta^{n}_{h_{k}}-\vartheta^{n}_{h_{k+1}}\|_{1}^{2}}.$
@@
To investigate spatial convergence, we compute the error between successive spatial refinements on meshes with $h_{k}\approx 2^{-k}$ for $k=4,\ldots,8$ while keeping the time step fixed at $\tau=2^{-10}\cdot 10^{-1}$. The final time is set to $T=10^{-1}$. The errors and approximate convergence rates, with $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, can be found in Table 1. They indicate optimal first-order convergence in the $H^{1}$-norm and the second-order convergence in the $L^{2}$-norm.
@@
4 $3.69\cdot 10^{-2}$ $9.15\cdot 10^{-3}$ $7.50\cdot 10^{-2}$ $1.97\cdot 10^{0}$
5 $\cdot 10^{-2} 0.820.82$ $5.63\cdot 10^{-3}$ $0.70$ $1.72\cdot 10^{-2}$ $2.12$ $9.92\cdot 10^{-1}$ $0.99$
6 $1.12\cdot 10^{-2}$ $0.90$ $2.69\cdot 10^{-3}$ $1.06$ $4.36\cdot 10^{-3}$ $1.98$ $5.04\cdot 10^{-1}$ $0.98$
7 $5.66\cdot 10^{-3}$ $0.98$ $1.27\cdot 10^{-3}$ $1.08$ $1.12\cdot 10^{-3}$ $1.95$ $2.50\cdot 10^{-1}$ $1.01$
To investigate time convergence, we fix the mesh size at $h=2^{-7}$ and consider time step sizes $\tau_{k}=2^{-k}\cdot 10^{-1}$ for k=7,,11k=7,...,11. The final time remains $T=10^{-1}$. The spatial mesh is fixed with mesh size $h=2^{-7}$. The temporal errors are defined analogously:
$\text{err}(\nabla\phi,\tau_{k})$ $:=\max_{n\in\mathcal{I}_{\tau_{k}}}\|\phi^{n}_{h,\tau_{k}}-\phi^{n}_{h,\tau_{k+1}}\|_{1},\penalty 10000\ \text{err}(\nabla\mu,\tau_{k}):=\sqrt{\tau_{k}\sum_{n=1}^{n_{\tau_{k}}}\|\mu^{n}_{h,\tau_{k}}-\mu^{n}_{h,\tau_{k+1}}\|_{1}^{2}},$
$\text{err}(\vartheta,\tau_{k})$ $:=\max_{n\in\mathcal{I}_{\tau_{k}}}\|\vartheta^{n}_{h,\tau_{k}}-\vartheta^{n}_{h,\tau_{k+1}}\|_{0},\penalty 10000\ \text{err}(\nabla\vartheta,\tau_{k}):=\sqrt{\tau_{k}\sum_{n=1}^{n_{\tau_{k}}}\|\vartheta^{n}_{h,\tau_{k}}-\vartheta^{n}_{h,\tau_{k+1}}\|_{1}^{2}}.$
@@
for $a\in\{\nabla\phi,\nabla\mu,\vartheta,\nabla\vartheta\}$. The errors and rates, for $\mathbf{C}=10^{-4}\cdot\mathbf{I}$, can be found in Table 2. As expected, optimal first-order convergence in time is achieved.
@@
7 $4.56\cdot 10^{-4}$ $0.86$ $8.93\cdot 10^{-5}$ $0.97$ $2.11\cdot 10^{-4}$ $1.00$ $3.53\cdot 10^{-3}$ $1.01$
8 $2.43\cdot 10^{-4}$ $0.91$ $4.52\cdot 10^{-5}$ $0.98$ $5\cdot 10^{-4} 1.001$.00 $1.76\cdot 10^{-3}$ $1.00$
9 $1.26\cdot 10^{-4}$ $0.94$ $2.27\cdot 10^{-5}$ $0.99$ $5.25\cdot 10^{-5}$ $1.00$ $8.81\cdot 10^{-4}$ $1.00$
10 $6.44\cdot 10^{-5}$ $0.97$ $1.14\cdot 10^{-5}$ $0.99$ $2.63\cdot 10^{-5}$ $1.00$ $4.40\cdot 10^{-4}$ $1.00$
@@ 6.2 Illustrating example
Figure 3: Snapshots of temperature $\vartheta$ for different cross-coupling matrices $\mathbf{C}$: (Top) $\mathbf{C}=0\cdot\mathbf{I}$; (Bottom) $\mathbf{C}=10^{-4}\cdot\mathbf{I}$.
@@ 7 Conclusion & Outlook
In this work, we proposed a fully discrete finite element method for the non-isothermal Cahn–Hilliard equation, based on a formulation using the entropy equation in place of the internal energy equation. The resulting scheme ensures mass conservation and entropy production, while the total energy—coinciding with the internal energy in this setting—is dissipated over time due to numerical diffusion. Our numerical experiments demonstrate optimal convergence rates: first-order in time and either first- or second-order in space, depending on the norm used. We also presented numerical results that highlight the influence of cross-coupling terms between the Cahn–Hilliard and entropy equations. Future work will aim at removing the artificial numerical dissipation to achieve discrete energy conservation. Additionally, we plan to extend the scheme to incorporate incompressible flow dynamics.
Acknowledgment
The present research has been supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) in the framework to the collaborative research center ”Multiscale Simulation Methods for Soft-Matter Systems” (TRR 146) under Project No. 233630050 and by the SPP 2256 ”Variational Methods for Predicting Complex Phenomena in Engineering Structures and Materials” under Project No. 441153493. M.L.-M. gratefully acknowledges the support of the Gutenberg Research Fellowship and the Mainz Institute for Multiscale Modelling (M3odel).